OpenStudy (anonymous):

If 9x - 13 = -31, find the value of x - 8

jimthompson5910 (jim_thompson5910):

What do you get when you solve 9x - 13 = -31 for x?

OpenStudy (anonymous):

x = -2

OpenStudy (anonymous):

Right? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

so x-8 is equal to?

OpenStudy (anonymous):

What does it mean when it says find the value of x - 8?

jimthompson5910 (jim_thompson5910):

well you know that x = -2, so replace x with -2 to get x - 8 = -2 - 8 = -10

jimthompson5910 (jim_thompson5910):

the value of x-8 is -10 (since x = -2)

jimthompson5910 (jim_thompson5910):

x is just a number (a place holder for a number) so is x-8 because subtracting 2 numbers results in another number

OpenStudy (anonymous):

I got -2 = -2 ?

OpenStudy (kl0723):

@gold_grace123 so if your answer is x=-2, then you'll have (-2)-8=-10 :)

OpenStudy (anonymous):

Alright, thanks guys!

jimthompson5910 (jim_thompson5910):

yeah exactly what I said